OpenStudy (anonymous):

can you help with this one? Simplify the number using the imaginary unit i. SQUARE ROOT -16

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

\(\sqrt{-16}\) \(\sqrt{-1 * 16}\) \(4*\sqrt{-1}\)

ganeshie8 (ganeshie8):

now use this :- \(i = \sqrt{-1}\) so it simplifies to \(4i\)

OpenStudy (anonymous):

thank you!!